OpenStudy (anonymous):

Can anyone give me the proper equation for this ??? you deposited $600 dollars in a savings account that earns a simple interest rate of 3.5% per year. How many years do you need to keep the money in the account to earn $210 in interest ?

OpenStudy (anonymous):

210 = 600*r*t 4=3.5/100